OpenStudy (thcsbabygirl1224):

Simplify the expression: 1/1+cot^2x sec2x csc2x sin2x cos2x tan2x

OpenStudy (thcsbabygirl1224):

@AloneS.

OpenStudy (freckles):

I guess the problem is 1/(1+cot^2(x)) ?

OpenStudy (freckles):

or in latex \[\frac{1}{1+\cot^2(x)}\] use a Pythagorean identity for the denominator

OpenStudy (thcsbabygirl1224):

i dont understand how you would use it im so lost

OpenStudy (freckles):

\[\sin^2(x)+\cos^2(x)=1 \\ \tan^2(x)+1=\sec^2(x) \\ 1+\cot^2(x)=\csc^2(x)\] these are the Pythagorean identities

OpenStudy (freckles):

you only need one of these here

OpenStudy (freckles):

can you make a guess to which one

OpenStudy (thcsbabygirl1224):

the last one because it has what your working with in it ?

OpenStudy (freckles):

right so are you saying you don't know how to use it here for this problem?

OpenStudy (thcsbabygirl1224):

no i tried and got something that was not there

OpenStudy (freckles):

what did you get

OpenStudy (thcsbabygirl1224):

cot^2 which idk how i got it tho

OpenStudy (freckles):

I don't know how you got that :p

OpenStudy (freckles):

you do know that by the Pythagorean identity we mentioned above that you can replace 1+cot^2(x) with csc^2(x) since 1+cot^2(x)=csc^2(x) ?

OpenStudy (thcsbabygirl1224):

?

OpenStudy (freckles):

do you see the 1+cot^2(x) on bottom do you know why I mentioned the Pythagorean identity 1+cot^2(x)=csc^2(x) it is because you have 1+cot^2(x) on bottom you can replace 1+cot^2(x) with csc^2(x) since 1+cot^2(x)= csc^2(x)

OpenStudy (alones.):

1 + tan^2(x) ___________ 1 + cot^2(x) 1 + tan^2(x) _____________ 1 + (1 / tan^2(x)) 1 + tan^2(x) ___________________ (1 + tan^2(x))/tan^2(x) = tan^2(x) b) sec^2(x) cot^2(x) = (1/cos^2(x))(cos^2(x)/sin^2(x)) = 1/sin^2(x) = csc^2(x)

OpenStudy (alones.):

This is what i got.

OpenStudy (thcsbabygirl1224):

so its d then since i finally think i got the work right

OpenStudy (freckles):

can I see what you think we have after using the Pythagorean identity I mentioned don't do anything else just use the Pythagorean identity I mentioned above

OpenStudy (thcsbabygirl1224):

brb class switch

OpenStudy (alones.):

k